what is vapour density of NH3?

Grade:12

5 Answers

sanjay
8 years ago
8.5
PRAPULPODISHETTI
8 years ago
682 kg/m is the density of liquid nh3
srujanyadav
8 years ago
682 kg/m3
L GOUSE BASHA
8 years ago
Ammonia Vapour Density (1.013 bar at boiling point) is 682 kg/m3
Beenu Mathew
7 years ago
Hi,
vapour density × 2 = molecular mass
NH3 = 14 + 3 = 17g
therefore
vapor density = 17/2 = 8.5g
CHEERS
HOPE THIS HELPS
GOOD LUCK!!